OpenStudy (ishipdestiel):

WILL FAN AND MEDAL FOR EXPLAINATION Using the following equation, find the center and radius of the circle. x^2 + 2x + y^2 + 4y = 20

OpenStudy (dayakar):

do u have any idea

OpenStudy (mathmale):

Caution: Please use the symbol ^ to express exponentiation: x2 + 2x + y2 + 4y = 20 NO x^2 + 2x + y^2 + 4y = 20 YES The standard equation of a circle centered at the point (h,k) and with radius r is:\[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (dayakar):

\[(x-h)^{2}+ (y - k)^{2}= r ^{2}\] is the standard form of circle equation whose center = (h ,k) and radius =r

OpenStudy (mathmale):

So, to rewrite x^2 + 2x + y^2 + 4y = 20 in the proper form (standard equation of a circle), you're going to have to "complete the square" twice in x^2 + 2x + y^2 + 4y = 20. Are you familiar with "completing the square?"

OpenStudy (ishipdestiel):

I have no idea what completeing the square means

OpenStudy (mathmale):

Let's look at the first two terms of x^2 + 2x + y^2 + 4y = 20. We want to re-write x^2 + 2x as the "perfect square" of a binomial. Right off, I'll share the correct end result: x^2 + 2x + 1 - 1. Please note that the first 3 terms are the "perfect square) of (x+1)." How to do this? Starting with x^2 + 2x, 1) Take ONE HALF of the coefficient of x. That's 2, so ONE HALF of that coeff. is 1. 2) Add 1, and then subtract 1, from x^2+2x. Result: x^2 + 2x + 1 - 1 = (x+1)^2 - 1 Any questions about this procedure?

OpenStudy (dayakar):

[x^2 +2*x*1 + 1^2]+[y^2 + 2 *y *2 +2^2] = 15 we are making the equation into (a+b) ^2 form

OpenStudy (dayakar):

(x+1)^2 + (y +2)^2 =sqrt 15

OpenStudy (dayakar):

\[(x+1)^{2}+ (y+2)^{2}= (\sqrt{15})^{2}\]

OpenStudy (ishipdestiel):

ok so why would you take Take ONE HALF of the coefficient of x. and then I understand from there

OpenStudy (dayakar):

\[a ^{2} + 2\times a \times b +b ^{2} = (a+b)^{2}\]compare \[x ^{2} + 2 \times x \times 1 + 1^{2} - 1^{2}\]

OpenStudy (dayakar):

do u understand

OpenStudy (ishipdestiel):

Not really where did these numbers come from

OpenStudy (mathmale):

"completing the square" is a commonly used method of algebra used for solving quadratic equations, locating the centers and radii of circles, and so on. Have you not found any mention of "completing the square" in your online learning materials or in your textbook?

OpenStudy (dayakar):

x^2 +2x +y^2 +4y + 5 = 15 can i write ur equation like this

OpenStudy (mathmale):

Let me ask the same questions then, in connection with your geometry course. Have you come across any example problems in which someone explains how to find the center and radius of a circle from a given equation such as x^2 + 2x + y^2 + 4y = 20 ? Aren't there any examples in your textbook or online learning materials?

OpenStudy (ishipdestiel):

I think, let me find it real quick

OpenStudy (mathmale):

Let's look at the "y-terms" of the given equation x^2 + 2x + y^2 + 4y = 20. First, let's go thru the procedure of completing the square applied to y^2 + 4y. Look at the coefficient of y here; it is 4. Take half of that: result is 2 Square this last result: next result is 4 Now add 4 and then subtract 4 from y^2+4y. Result: y^2+4y + 4 - 4. Note that y^2+4y+4 is a "perfect square," equal to (y+2)^2. Thus, your final result would be (y+2)^2-4. You need to understand the various steps of this procedure. With a little practice, the procedure becomes almost automatic, easy to do. Can you go back to x^2 + 2x and complete the square, following the same steps shown immediately above?

OpenStudy (mathmale):

I see you've tried to share something from FLVS, but since I don't have the password, I can't access it. Sorry. Either find another way to share this info (as thru sharing a screen shot), or let's go thru my examples again. Your choice.

OpenStudy (mathmale):

Let's go thru this problem again, with your participation: Given: x^2 + 2x + y^2 + 4y = 20 Let's "complete the square" for x^2 + 2x. What is the coefficient of "x" in x^2 + 2x? (Awaiting your answer)

OpenStudy (ishipdestiel):

The coefficient of x i think would be 2, then I half it and get 1 and 1 squared= 1 After that I dont understand the " Now add 4 and then subtract 4 from y^2+4y." from your example

OpenStudy (mathmale):

Your given equation has two parts on the left side: x terms and y terms. You are correct: x^2 + 2x becomes x^2 + 2x + 1 - 1, or (x+1)^2 -1. Use exactly the same steps to rewrite y^2 + 4y. Awaiting your answer.

OpenStudy (ishipdestiel):

ok so the coefficint of y would be 4. I would then half it, getting 2, then i square that and get 4,so (y+2)^2-4, which for both of them put together would be (x+1)^2 -1+ (y+2)^2-4= 20

OpenStudy (bobo-i-bo):

yup, now put the -1 and the -4 on the other side and you should recognise that the equation is the normal form of the equation for a circle.

OpenStudy (mathmale):

So, please add 1+4 (or 5) to both sides of your latest equation. This will remove the -4 and the -1 from the left side. What do you get? You'll most likely have the desired equation of this circle in standard form. What's 20 + 1 + 4? If that's a perfect square, find the square root (r) and rewrite the right side as r^2.